OpenStudy (anonymous):

So this one is almost exactly like the last one but I'm still doing something wrong

OpenStudy (anonymous):

\[\int\limits_{}^{}x^4\sqrt{5+x^5}dx\] u=\[5+x^{5}\] \[\frac{ du }{ dx }=5x ^{4}\] du=5x^4 dx

OpenStudy (anonymous):

yes \[ I=\int\sqrt{5+x^5}(x^4dx)=\int\sqrt{u}\frac{du}{5} \]

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }du=x^{4}dx\] \[\frac{ 1 }{ 5 } \int\limits_{}^{}\sqrt{u} du\]

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

so why isn't webworks accepting it?

OpenStudy (anonymous):

show your final answer

OpenStudy (anonymous):

(sqrt(5+x^5))/5+C

OpenStudy (anonymous):

see

OpenStudy (anonymous):

\[ I={1\over5}\int u^{1\over2}du={1\over5}\frac{u^{3\over2}}{3\over2}+C\\ I={1\over5}{2\over3}u\sqrt{u}+C\\ I={2\over15}(1+x^5)\sqrt{1+x^5}+C \]

OpenStudy (anonymous):

now I'm lost again

OpenStudy (anonymous):

the first statement is the one right after the substitution, right?

OpenStudy (anonymous):

i think you did everything but skipped the actual "integration": part

OpenStudy (anonymous):

so I get that you changed the \[\sqrt{u} \to u^{1/2}\] and then took the antiderivative of that

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

which is \[\frac{ 1 }{ 5 }*\frac{ 2 }{ 3 }u ^{3/2}+C\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

but then why/how did you go back to \[\sqrt{u}\]

OpenStudy (anonymous):

ooh \[u^{3/2}=u\times u^{1/2}=u\sqrt{u}\]

OpenStudy (anonymous):

I do not like exponent.. this way looks waaay cooler

OpenStudy (anonymous):

ok so then keeping it my way would be \[\frac{ 1 }{ 5 }\frac{ 2 }{ 3 }(5+x^5)^{2/3}+C\]

OpenStudy (anonymous):

yes . perfectly fine

OpenStudy (anonymous):

which is sloppy, but I get it and webworks accepts it

OpenStudy (agent0smith):

@heradog in your last post you have the exponent as 2/3 not 3/2.

OpenStudy (anonymous):

oops, I had it right everywhere else, thanks